# Entropic repulsion and scaling limit for a finite number of non-intersecting subcritical FK interfaces

### v1

@@ 1 Introduction
Of the same nature is the study of a system of multiple interfaces interacting together. Such a problem will be the focus of this work. Indeed, this paper determines the scaling behaviour of a finite number of long clusters of subcritical Fortuin–Kasteleyn (FK) percolation, conditioned not to intersect. Regarding the discussion above, it can be seen as the study of a finite number of interfaces interacting together. An interesting feature of this setting is that when conditioned on not intersecting, the interaction between the clusters can turn out to be attractive, a priori allowing the existence of a pinning transition — a regime where this attraction is so strong that the clusters actually remain at a bounded distance from each other.
@@
In this paper, we rule out the existence of such a transition. In the fashion of [ottvelenikwachtelioffe], we show that the behaviour of this system obeys an entropic repulsion phenomenon: the entropy caused by the large number of possible interfaces wins over the energetic reward obtained by staying close to each other, all the way down to the critical temperature. Such a phenomenon has been previously identified in a variety of settings, for instance in the three-dimensional semi-infinite Ising model at low temperatures [frohlichpfisterSemiInfiniteIsing], the 2+1-dimensional SOS model above a hard wall [caputomartinellitoninelliEntropicRepulsionSOS], a 1+1-dimensional interface above an attractive field in presence of a magnetic field [Ve2004] or a supercritical Potts interface above a wall [ottvelenikwachtelioffe], to mention but a few works studying this phenomenon.
@@
In this work, entropic repulsion of the FK clusters at any subcritical temperature is established in Proposition LABEL:proposition_global_entropic_repulsion, which is probably the most important output of this work. As a byproduct, we derive two results regarding the global behaviour of such a system of conditioned interfaces. The first one is the diffusive scaling limit of such a system, which is shown to be a system of Brownian bridges conditioned not to intersect: the so-called Brownian watermelon. Moreover, we observe that the entropic repulsion phenomenon also allows the computation — up to a multiplicative constant — of the probability of the existence of such a system of interfaces. Finally, as a byproduct of the latter observation, we also obtain the asymptotics of the probability of the occurrence of a large finite connection in the supercritical random-cluster model.
@@
The method is in spirit close to that of [ottvelenikwachtelioffe], but we face some additional difficulties — essentially due to the fact that the interaction is not only between a random interface and a deterministic object as in the latter work, but between several random interfaces, making the proofs more intricate as one has to control the joint behaviour of these interacting objects: in particular Section LABEL:section_RCM contains some estimates concerning the simultaneous behaviour of the system of clusters that appear to be new. The proofs crucially rely on the Ornstein–Zernike theory for subcritical random-cluster models, developed in [ozrandomclustercampanino].
@@ 1.1.1 The random-cluster model
$\partial G=\left\{x\in V(G),\exists y\notin V(G),\{x,y\}\in E(\mathbb{Z}^{2})\right\}\\ (\text{resp. }\partial_{\mathsf{ext}}G=\left\{x\notin V(G),\exists y\in V(G),\{x,y\}\in E(\mathbb{Z}^{2})\right\}).$

### v2

@@ 1 Introduction
Of the same nature is the study of a system of multiple interacting interfaces, which is the focus of the present work. Indeed, this paper determines the scaling behaviour of a finite number of long clusters of subcritical Fortuin–Kasteleyn (FK) percolation, conditioned not to intersect; subcritical percolation clusters mimic interfaces in the low temperature regime.
@@
An interesting feature of this setting is that when conditioned not to intersect, the interaction between the clusters can turn out to be attractive, a priori allowing the existence of a pinning transition — a regime where this attraction is so strong that the clusters actually remain at a bounded distance from each other. We rule out the existence of such a transition. In the fashion of [ottvelenikwachtelioffe], we show that the behaviour of this system obeys an entropic repulsion phenomenon: the entropy caused by the large number of possible clusters wins over the energetic reward obtained by keeping them close together, all the way up to the critical point. Such a phenomenon has been previously identified in a variety of settings, for instance in the three-dimensional semi-infinite Ising model at low temperatures [frohlichpfisterSemiInfiniteIsing], the 2+1-dimensional SOS model above a hard wall [caputomartinellitoninelliEntropicRepulsionSOS], a 1+1-dimensional interface above an attractive field in presence of a magnetic field [Ve2004] or a supercritical Potts interface above a wall [ottvelenikwachtelioffe], to mention but a few works studying this phenomenon.
@@
In this work, entropic repulsion of the FK clusters at any subcritical temperature is established in Proposition LABEL:proposition_global_entropic_repulsion, which is probably the most important output of this work. As a byproduct, we derive two results regarding the global behaviour of such a system of conditioned clusters. The first one is the diffusive scaling limit of such a system, which is shown to be a system of Brownian bridges conditioned not to intersect: the so-called Brownian watermelon. Moreover, we observe that the entropic repulsion phenomenon also allows the computation — up to a multiplicative constant — of the probability of the existence of such a system of clusters. Finally, as a byproduct of the latter observation, we also obtain the asymptotics of the probability of the occurrence of a large finite connection in the supercritical random-cluster model.
@@
The method is in spirit close to that of [ottvelenikwachtelioffe], but with considerable additional difficulties. These are essentially due to the fact that the interaction is not only between a random object and a deterministic one, but between several random objects, forcing one to control their joint behaviour. The proofs make heavy use of the Ornstein–Zernike theory for subcritical random-cluster models, developed in [ozrandomclustercampanino].
@@ 1.1.1 The random-cluster model
$\partial G$ $=\left\{x\in V(G),\exists y\notin V(G),\{x,y\}\in E(\mathbb{Z}^{2})\right\}\text{ and}$
$\partial_{\mathsf{ext}}G$ $=\left\{x\notin V(G),\exists y\in V(G),\{x,y\}\in E(\mathbb{Z}^{2})\right\},\text{ respectively.}$